Express each percent as a fraction in simplest form.
91 2/3%

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

91 2/3= [(3*91)+2]/3=(273+2)/3=275/3

Answer by stanbon(29603)

Express each percent as a fraction in simplest form.
91 2/3%
=(275/3)%
=(275/[3*(100/3)]
= [27500/3][1/100]
= [27500/3]%
= 9166.667%
==============
Another way to do this:
91 2/3%
91.666666666666
= 9166.66666..../100
= 9166.6666...%
=====================
Cheers,
Stan H.
Cheers,
Stan H.